Can you please help me answer questions a, b, c, d, and e below? Thanks.Problem 20-29 (LO. 1, 4) Pink Corporation acquired land and securities in a $ 351 tax-free exchange. On the date of the transC. Assume there is no business reason for the transfer of the property. If the land is distributed 85% to Maria and 15% to Pa
Answer #1

a. realized loss = 500,000 - 720,000 = 220,000

b. Built in loss = 90,000 [ 630,000 - 720,000]

c. Loss on distribution = 220,000 *85% = 187,000

Built in loss = 90,000 *15% = 13500

Loss recognized = 220,000 - 187,000 - 13500 = 19500

d. Loss on distribution to Maria = 110,000 [ 220,000*50%]

Loss on distribution to Paul = 110,000

Loss disallowed = 110,000

e. Built in loss =$90,000

Any doubt please comment.